An investor sold her stock in two companies, A and B, for $44,200 and $2,800, respectively. Each stock had been bought for an initial purchace of $1000, with all dividends reinvested. The A stock had been held for 13.26 years, while the B stock had been held for 5.75 years. For each stock, compute the compound interest rate which would yield an equivalent return over the same time period. Which was the more lucrative investement?
Answer #1

Compound interest rate on the stock sold to to company A is calculated below:

1000 X (1 + i)13.26 = 44,200

(1 + i)13.26 = 44.2

(1+i) = 44.213.26

(1+i) = 1.3307

i = 0.3307

Compound interest rate on the stock sold to to company B is calculated below:

1000 X (1+i)5.75 = 2,800

(1+i) 5.75 = 2.8

(1 + i) = 2.85.75

(1+i) = 1.1961

i = 0.1961

Investment in company A is more lucrative as it gives higher compound interest rate.